OpenStudy (anonymous):

Which of the following points is a solution to the system of equations? x + 2y = 11 3x - 4y = -7 Question 6 options: (3,4) (-7,4) (-1,6) (3,7)

OpenStudy (anonymous):

(3) + 2(4) = 3+8 =11 3(3) - 4(4) = 9-16= -7